OpenStudy (anonymous):

If g(x) = x2 + 2, find g(3).

OpenStudy (zzr0ck3r):

x^2?

OpenStudy (zzr0ck3r):

g(3) = (3)^2 + 2 = ?

OpenStudy (anonymous):

yes, sorry

OpenStudy (zzr0ck3r):

what is 3^2?

OpenStudy (zzr0ck3r):

3^2 = 3*3

OpenStudy (anonymous):

9

OpenStudy (zzr0ck3r):

what is 9 + 2?

OpenStudy (anonymous):

11

OpenStudy (anonymous):

good job

OpenStudy (anonymous):

what is a function?

OpenStudy (zzr0ck3r):

lol the definition of a function is a little outside of the range of what he is doing:)

OpenStudy (anonymous):

look at you ZZ lol clever eh

OpenStudy (zzr0ck3r):

ha:)

OpenStudy (anonymous):

bella, check out pp 39 and on in this website: http://finedrafts.com/files/Larson%20PreCal%208th/Larson%20Precal%20CH1.pdf